A triangle has angle measures of 2x+10, 4x and 5x+5. What are the measures of each angle from smallest to largest?
Answer by nycsharkman(136)

A triangle has angle measures of 2x+10, 4x and 5x+5. What are the measures of each angle from smallest to largest?
All triangles measure 180 degrees.
You were given 3 angles.
Add all 3 angles, equate sum to 180 and solve for x.
2x + 10 + 4x + 5x + 5 = 180
11x + 15 = 180
11x = 180 - 15
11x = 165
x = 165/11
x = 15
We just found x to be 15, right?
The angles are 2x + 10, 4x and 5x + 5.
Replace x with 15 and do the math.
2(15) + 10 = 30 + 10 = 40
Next: 4(15) = 60
Next: 5(15) + 5 = 75 + 5 = 80
From smallest to largest:
40, 60, 80